# Adapting Differentially Private Synthetic Data to Relational Databases

### v1

@@ Appendix B More Details about Algorithm 1
     Solve $\bm{b}^{\text{syn}}={\arg\min}_{\begin{subarray}{c}\bm{b}\in[0,1]^{n_{1}^{\text{syn}}\cdot n_{2}^{\text{syn}}}\\ \bm{1}^{T}\bm{b}=m^{\text{syn}}\end{subarray}}\lVert\frac{1}{m^{\text{syn}}}\hat{\bm{Q}}\bm{b}-\hat{\bm{a}}\rVert_{2}^{2}$ using mirror $\triangleright$ Algorithm 5
     $\bm{B}^{\text{syn}}=\operatorname{RandomizedRounding}(\bm{b}^{\text{syn}}.\text{reshape}(n_{1}^{\text{syn}},n_{2}^{\text{syn}}))$$\triangleright$ Algorithm 6
@@
Assert $\operatorname{sum}(\bm{B})==m^{\text{syn}}$
Initialize: $\mathcal{O}=\emptyset$ and $\mathcal{W}=\mathsf{Subsample}\left(\{(\mathcal{S}_{1}^{1},\mathcal{S}_{2}^{1}),\cdots,(\mathcal{S}_{1}^{W},\mathcal{S}_{2}^{W})\}\backslash\mathcal{T}_{\text{selected}}\right)$$\triangleright$ Save running time
@@
     $\mathsf{score}(w)=\frac{1}{2}\|P_{w}(\mathcal{B})-P_{w}(\mathcal{B}^{\text{syn}})\|_{1}$$\triangleright$ total variance distance between (empirical) marginal distributions of $\mathcal{B}$ and $\mathcal{B}^{\text{syn}}$ on workload $w$
@@
Algorithm 5 Using mirror descent for solving bounded-variable least-squares problem.
Input: Observed query matrix $\bm{Q}$; noisy observations $\hat{\bm{a}}$; step size $\eta$; number of iterations for the outer and inner loop $T_{o},T_{i}$; initialization $\bm{b}^{0}$; number of relationship in synthetic tables $m^{\text{syn}}$
Initialize: $\bm{b}=\bm{b}^{0}/m$, $N=\operatorname{dim}(\bm{b})$
for $t=1,\cdots,T_{o}$ do
     $\bm{g}=\nabla\|\frac{1}{m^{\text{syn}}}\hat{\bm{Q}}\bm{b}-\hat{\bm{a}}\|_{2}^{2}=2\frac{1}{m^{\text{syn}}}\bm{Q}^{T}(\frac{1}{m^{\text{syn}}}\bm{Q}\bm{b}-\hat{\bm{a}})$
     $b_{i}=m^{\text{syn}}\cdot\frac{b_{i}\operatorname{exp}(-\eta g_{i})}{\sum_{j=1}^{N}b_{j}\operatorname{exp}(-\eta g_{j})},\quad\forall i\in[N]$
     for $l=1,\cdots,T_{i}$ do
         $\bm{b}=\min(\bm{b},\bm{1})$
         $b_{i}=m^{\text{syn}}\cdot\frac{b_{i}}{\sum_{j=1}^{N}b_{j}},\quad\forall i\in[N]$
@@
end for
Output: $\bm{b}^{\text{syn}}=\bm{b}$
Algorithm 6 Randomized rounding for many-to-many databases.
Input: relaxed bi-adjacency matrix $\bm{B}\in[0,1]^{n_{1}^{\text{syn}}\times n_{2}^{\text{syn}}}$, number of relationship in synthetic tables $m^{\text{syn}}$
Assert $\operatorname{sum}(\bm{B})==m^{\text{syn}}$
Initialize: $m=0$, $\mathcal{E}=\emptyset$, $\bm{B}^{\text{syn}}=\bm{0}^{n_{1}^{\text{syn}}\times n_{2}^{\text{syn}}}$
while $m<m^{\text{syn}}$ do
     Draw a random sample $(x_{i},x_{j})$ from $[n_{1}^{\text{syn}}]\times[n_{2}^{\text{syn}}]$ with probability $\Pr(i,j)=\frac{B_{i,j}}{m^{\text{syn}}}$
     if $(x_{i},x_{j})\notin{\mathcal{E}}$ then
         =${\mathcal{E}}={\mathcal{E}}\cup\{(x_{i},x_{j})\},\quad m=m+1$
         $B^{\text{syn}}_{x_{i},x_{j}}=1$
     end if
end while
Output: $\bm{B}^{\text{syn}}$
Appendix C Omitted Proofs
We start this section by proving the DP guarantee for our main algorithm. We recall Theorem 1 below for reference.
Theorem 3.
Suppose we use any DP mechanism with privacy budgets $(\varepsilon_{1},\delta_{1})$ and $(\varepsilon_{2},\delta_{2})$ to generate individual synthetic tables, respectively. Then we apply Algorithm  1 with privacy budget $(\varepsilon_{\text{rel}},\delta_{\text{rel}})$ to build their relationship. The synthetic relational database satisfies $(\varepsilon_{1}+\varepsilon_{2}+\varepsilon_{\text{rel}},\delta_{1}+\delta_{2}+\delta_{\text{rel}})$-DP, where $\varepsilon_{\text{rel}}(\delta_{\text{rel}})\leq\rho_{\text{rel}}+2\sqrt{\rho_{\text{rel}}\log(1/\delta_{\text{rel}})}$. This condition can be achieved by choosing $\rho_{\text{rel}}=\left(\sqrt{\log(1/\delta_{\text{rel}})+\epsilon_{\text{rel}}}-\sqrt{\log(1/\delta_{\text{rel}})}\right)^{2}$.
Proof.
At each iteration, Algorithm  1 makes $K$ call to the exponential mechanism and $K$ call to the Gaussian mechanism. The exponential mechanism uses the score function:
@@ Appendix C Omitted Proofs
Recall that we consider two relational databases $\mathcal{B}$ and $\mathcal{B}^{\prime}$ adjacent if $\mathcal{B}^{\prime}$ can be obtained from $\mathcal{B}$ by selecting a table, modifying the values of a single row within this table, and changing all relationship associated with this row. We focus on the setting of bounded DP so both $\mathcal{B}$ and $\mathcal{B}^{\prime}$ have $m$ relationships. Additionally, we assume each record can have a relationship with at most $d_{\max}$ records in the other table. Hence, for any workload $w$, the total variance distance between the empirical distributions of $\mathcal{B}$ and $\mathcal{B}^{\prime}$ on $w$ has an upper bound:
$\frac{1}{2}\|P_{w}(\mathcal{B})-P_{w}(\mathcal{B}^{\prime})\|\leq\frac{d_{\max}}{m}.$ (10)

### v2

@@ B More Details about Algorithm 1
    Randomly choose $\bm{v}^{\text{slice}}_{1}\in\{0,1\}^{n_{1}^{\text{syn}}}$, $\bm{v}^{\text{slice}}_{2}\in\{0,1\}^{n_{2}^{\text{syn}}}$ s.t. $\mathsf{sum}({\bm{v}^{\text{slice}}_{i}})=\lfloor\tau n_{i}^{\text{syn}}\rfloor$ for $i\in\{1,2\}$
    Set $\bm{B}^{\text{syn}}_{\text{slice}}=[\bm{v}^{\text{slice}}_{1}]\bm{B}^{\text{syn}}[\bm{v}^{\text{slice}}_{2}]$ and $\bm{Q}_{\text{slice}}=\bm{Q}[\operatorname{vec}(\bm{v}^{\text{slice}}_{1}\cdot(\bm{v}^{\text{slice}}_{2}.\text{transpose}))]$$\triangleright$ Definition 5
    Compute $m_{\text{slice}}=\mathsf{sum}(\bm{B}^{\text{syn}}_{\text{slice}})$
    Solve the relaxed optimization:$\triangleright$ Algorithm 2
$\widetilde{\bm{B}}^{\text{syn}}_{\text{slice}}=\operatorname*{\arg\!\min}_{\begin{subarray}{c}B_{i,j}\in[0,1]\\ \mathsf{sum}(\bm{B}_{\text{slice}}^{\text{syn}})=m_{\text{slice}}\end{subarray}}\lVert\frac{1}{m_{\text{slice}}}\bm{Q}_{\text{slice}}\mathsf{flatten}(\bm{B}_{\text{slice}}^{\text{syn}})-\hat{\bm{a}}\rVert_{2}^{2}$
    $\bm{B}^{\text{syn}}_{\text{slice}}=\operatorname{UnbiasedSampling}(\mathsf{flatten}(\widetilde{\bm{B}}^{\text{syn}}_{\text{slice}}),m_{\text{slice}})$$\triangleright$ Algorithm 3
    $[\bm{v}^{\text{slice}}_{1}]\bm{B}^{\text{syn}}[\bm{v}^{\text{slice}}_{2}]=\bm{B}^{\text{syn}}_{\text{slice}}$$\triangleright$ Re-insert $\bm{B}^{\text{syn}}_{\text{slice}}$ into $\bm{B}^{\text{syn}}$
@@
Initialize: $\mathcal{O}=\emptyset$ and $\mathcal{W}=\mathsf{Subsample}\left(\{(\mathcal{S}_{1}^{1},\mathcal{S}_{2}^{1}),\cdots,(\mathcal{S}_{1}^{W},\mathcal{S}_{2}^{W})\}\backslash\mathcal{T}_{\text{selected}}\right)$$\triangleright$ Save runtime
@@
    $\mathsf{score}(w)=\frac{1}{2}\|P_{w}(\mathcal{B})-P_{w}(\mathcal{B}^{\text{syn}})\|_{1}$$\triangleright$ total variance distance between marginal distributions of $\mathcal{B}$ and $\mathcal{B}^{\text{syn}}$ on workload $w$
@@
C One-to-Many Relationship
In the main body of the paper, we focused on many-to-many relationship. In some cases, the original database may exhibit a one-to-many relationship. For instance, in the context of education data (e.g., student and school data), each school can have multiple students, but each student belongs to only one school. When generating a synthetic database, it is crucial to preserve this relationship. In this section, we apply our algorithm to this scenario. As shown, by incorporating this additional assumption—that records in one table have exactly one relationship with the other table—our algorithm becomes significantly simpler and operates much more efficiently.
Consider that each record in the first table of the original database has only one corresponding relationship (e.g., the student table from the example above). We denote the rows of the bi-adjacency matrix $\bm{B}^{\text{syn}}$ as $\bm{b}_{1},\cdots,\bm{b}_{n_{1}^{\text{syn}}}$. We introduce the constraint $\bm{1}^{T}\bm{b}_{i}=1$ for all $i\in[n_{1}^{\text{syn}}]$. After applying the unbiased sampling algorithm, this constraint will translate into the condition that each record in the first synthetic table has exactly one relationship with a record in the second table. Finally, we denote the concatenation of $\bm{b}_{1},\cdots,\bm{b}_{n_{1}^{\text{syn}}}$ by $\bm{b}$ (i.e., $\bm{b}$ is the vectorization of $\bm{B}^{\text{syn}}$). We solve the following relaxed optimization:
$\min_{\bm{b}\in\Omega}~{}f(\bm{b})\quad\text{where }$ $\ f(\bm{b})\triangleq\left\|\frac{1}{m^{\text{syn}}}\hat{\bm{Q}}\bm{b}-\hat{\bm{a}}\right\|_{2}^{2}$ (14)
and $\ \Omega\triangleq\{\bm{b}\mid 0\leq\bm{b}\leq 1\text{ and }\bm{1}^{T}\bm{b}_{i}=1\ \forall i\in[n_{1}^{\text{syn}}]\}.$
Note that the constraint set $\Omega$ can be reformulated for each individual vector $\bm{b}_{i}$:
$\Omega_{i}\triangleq\{\bm{b}_{i}\mid 0\leq\bm{b}_{i}\leq 1,\bm{1}^{T}\bm{b}_{i}=1\}.$
Therefore, we can still apply the projected gradient descent algorithm to solve the relaxed optimization problem in (14). In each iteration, after computing the gradient (8) based on the last iterate, we update the solution in the direction opposite to the gradient. Instead of projecting the entire vector $\bm{b}$ onto $\Omega$, we project each sub-vector $\bm{b}_{i}$ onto its respective constraint set $\Omega_{i}$. This can be achieved by applying the same closed-form expression from Theorem 2 to each $\bm{b}_{i}$. The projection step maintains the same time complexity but offers the added advantage of allowing parallel projection of the sub-vectors onto their respective constraint sets. Lastly, since the objective function remains unchanged, the convergence guarantee and learning rate are preserved.
Our projected gradient descent algorithm ensures that each row of $\widetilde{\bm{B}}^{\text{syn}}$ is a valid probability vector. Hence, we can directly sample from the categorical distribution parameterized by each row. We introduce Algorithm 7 that samples from $\widetilde{\bm{B}}^{\text{syn}}$ while ensuring the one-to-many relationship. This approach provides an unbiased sample and has the added benefit of enabling parallel sampling for each record in the first table.
Algorithm 7 Categorical rounding for one-to-many relationship.
Input: relaxed bi-adjacency matrix $\widetilde{\bm{B}}^{\text{syn}}\in[0,1]^{n^{\text{syn}}_{1}}\times[0,1]^{n^{\text{syn}}_{2}}$
Assert $\operatorname{sum}(\widetilde{\bm{B}}^{\text{syn}}_{i,:})==1$ for $i\in\{1,\cdots,n_{1}^{\text{syn}}\}$
$\bm{B}^{\text{syn}}=\mathsf{zeros}(n_{1}^{\text{syn}},n_{2}^{\text{syn}})$
for $i=1,\cdots,n_{1}^{\text{syn}}$ do
    Draw $k$ from $\{1,\cdots,n_{2}^{\text{syn}}\}$ according to the probability distribution $\widetilde{\bm{B}}^{\text{syn}}_{i,:}$
    $B_{i,k}^{\text{syn}}=1$
@@ C One-to-Many Relationship
Output: unweighted bi-adjacency matrix $\bm{B}^{\text{syn}}$
In short, generating synthetic data becomes significantly easier when the original relational database only contains one-to-many relationships. In our case, both the projected gradient descent algorithm and the unbiased sampling algorithm are simplified, enabling more efficient execution through parallel computation.
D Omitted Proofs
We provide omitted proofs in this section.
D.1 Proof of Lemma 1
Proof  Recall from Section 2.2 that for any $k$-way cross-table query $\bm{q}_{(\mathcal{S}_{1},\mathcal{S}_{2}),(\bm{y}_{1},\bm{y}_{2})}$, its average over a relational database $(\mathcal{D}_{1},\mathcal{D}_{2},\bm{B})$ can be computed as
$\bm{q}_{(\mathcal{S}_{1},\mathcal{S}_{2}),(\bm{y}_{1},\bm{y}_{2})}(\mathcal{B})=\frac{1}{\bm{1}^{T}\cdot\operatorname{vec}(\bm{B})}\sum_{i=1}^{n_{1}}\sum_{j=1}^{n_{2}}\bm{q}_{(\mathcal{S}_{1},\mathcal{S}_{2}),(\bm{y}_{1},\bm{y}_{2})}(\bm{x}_{i},\bm{x}_{j})B_{i,j}$ (15)
where $\bm{x}_{i}\in\mathcal{D}_{1}$ and $\bm{x}_{j}\in\mathcal{D}_{2}$. We define $\bm{Q}_{(\mathcal{S}_{1},\mathcal{S}_{2}),(\bm{y}_{1},\bm{y}_{2})}\triangleq\bm{1}_{(\mathcal{S}_{1},\bm{y}_{1})}\cdot\bm{1}^{T}_{(\mathcal{S}_{2},\bm{y}_{2})}\in\mathbb{R}^{n_{1}\times n_{2}}$ with its $i$-th row and $j$-th column denoted by $\left[\bm{Q}_{(\mathcal{S}_{1},\mathcal{S}_{2}),(\bm{y}_{1},\bm{y}_{2})}\right]_{i,j}$. By the definition of $\bm{1}_{(\mathcal{S}_{1},\bm{y}_{1})}$ and $\bm{1}_{(\mathcal{S}_{2},\bm{y}_{2})}$, we have $\left[\bm{Q}_{(\mathcal{S}_{1},\mathcal{S}_{2}),(\bm{y}_{1},\bm{y}_{2})}\right]_{i,j}=\bm{q}_{(\mathcal{S}_{1},\mathcal{S}_{2}),(\bm{y}_{1},\bm{y}_{2})}(\bm{x}_{i},\bm{x}_{j})$. Therefore,
$\operatorname{vec}\left(\bm{Q}_{(\mathcal{S}_{1},\mathcal{S}_{2}),(\bm{y}_{1},\bm{y}_{2})}\right)^{T}\cdot\operatorname{vec}(\bm{B})$ $=\sum_{i=1}^{n_{1}}\sum_{j=1}^{n_{2}}[\bm{Q}_{(\mathcal{S}_{1},\mathcal{S}_{2}),(\bm{y}_{1},\bm{y}_{2})}]_{i,j}B_{i,j}$
$=\sum_{i=1}^{n_{1}}\sum_{j=1}^{n_{2}}\bm{q}_{(\mathcal{S}_{1},\mathcal{S}_{2}),(\bm{y}_{1},\bm{y}_{2})}(\bm{x}_{i},\bm{x}_{j})B_{i,j}.$
Combined this with (15) yields the desired result.
D.2 Proof of Theorem 1
We recall Theorem 1 below and provide additional details along with the proof.
Theorem 7
Assume that each record in the original relational database has at most $d_{\max}$ relationships with records in the other table. We apply any DP mechanism with privacy budgets $(\epsilon_{1},\delta_{1})$ and $(\epsilon_{2},\delta_{2})$ to generate individual synthetic tables, respectively. Then we apply Algorithm  1 with privacy budget $(\epsilon_{\text{rel}},\delta_{\text{rel}})$ to build their relationship. The synthetic relational database satisfies $(\epsilon_{1}+\epsilon_{2}+\epsilon_{\text{rel}},\delta_{1}+\delta_{2}+\delta_{\text{rel}})$-DP, where $\epsilon_{\text{rel}}(\delta_{\text{rel}})\leq\rho_{\text{rel}}+2\sqrt{\rho_{\text{rel}}\log(1/\delta_{\text{rel}})}$. This condition can be achieved by choosing $\rho_{\text{rel}}=\left(\sqrt{\log(1/\delta_{\text{rel}})+\epsilon_{\text{rel}}}-\sqrt{\log(1/\delta_{\text{rel}})}\right)^{2}$.
Proof  At each iteration, Algorithm  1 makes $K$ call to the exponential mechanism and $K$ call to the Gaussian mechanism. The exponential mechanism uses the score function:
@@ D.2 Proof of Theorem 1
Recall that we consider two relational databases $\mathcal{B}$ and $\mathcal{B}^{\prime}$ adjacent if $\mathcal{B}^{\prime}$ can be obtained from $\mathcal{B}$ by selecting a table, modifying the values of a single row within this table, and changing all relationship associated with this row. We focus on the setting of bounded DP so both $\mathcal{B}$ and $\mathcal{B}^{\prime}$ have $m$ relationships. Additionally, we assume each record can have at most $d_{\max}$ relationships with records in the other table. Hence, for any workload $w$, the total variance distance between the empirical distributions of $\mathcal{B}$ and $\mathcal{B}^{\prime}$ on $w$ has an upper bound:
$\frac{1}{2}\|P_{w}(\mathcal{B})-P_{w}(\mathcal{B}^{\prime})\|_{1}\leq\frac{d_{\max}}{m}.$ (17)